# A Finite-Time Analysis of Distributed Q-Learning

### v1

@@ Abstract
Multi-agent reinforcement learning (MARL) has witnessed a remarkable surge in interest, fueled by the empirical success achieved in applications of single-agent reinforcement learning (RL). In this study, we consider a distributed Q-learning scenario, wherein a number of agents cooperatively solve a sequential decision making problem without access to the central reward function which is an average of the local rewards. In particular, we study finite-time analysis of a distributed Q-learning algorithm, and provide a new sample complexity result of $\tilde{\mathcal{O}}\left(\min\left\{\frac{1}{\epsilon^{2}}\frac{t_{\text{mix}}}{(1-\gamma)^{6}d_{\min}^{4}},\frac{1}{\epsilon}\frac{\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\sigma_{2}(\bm{W}))(1-\gamma)^{4}d_{\min}^{3}}\right\}\right)$ under tabular lookup setting for Markovian observation model.
@@ 1 Introduction
MARL has been studied under various scenarios including an access to central reward function [Tan, 1993, Claus and Boutilier, 1998, Littman, 2001, Mathkar and Borkar, 2016]. In particular, our interest lies in the the distributed learning paradigm where agents collaborate to solve a shared problem, constrained to communicate solely with their neighboring agents and does not have access to central reward function. Such setting has came of interest due to its wide applications [Zhao et al., 2021]. Compared to scenarios where a centralized coordinate exists, the distributed paradigm has advantage in terms of privacy-preservation and scalability. One notable example is the distributed adaptation of temporal-difference (TD) learning, as demonstrated in studies by Doan et al. [2019], Sun et al. [2020], Wang et al. [2020], Lim and Lee [2023], to name a few.
Meanwhile, in the literature of single-agent RL, Q-learning [Watkins and Dayan, 1992] is one of the most important algorithms in RL. The non-linear max-operator in Q-learning algorithm imposes difficulty in the analysis, and its non-asymptotic analysis has been an active research area recently [Even-Dar et al., 2003, Chen et al., 2021, Lee et al., 2023, Li et al., 2024]. However, distributed learning framework for Q-learning has not been studied in detail. In particular, distributed Q-learning has been studied in an asymptotic sense [Kar et al., 2013], i.e., the algorithm converges over time as it approaches infinity, or in a non-asymptotic sense under additional assumptions on the problem [Heredia et al., 2020, Zeng et al., 2022b]. This motivates our study to understand its non-asymptotic behavior under tabular setup, i.e., all the state-action values are stored in a table. Our contribution can be summarized as follows:
@@
    For Markovian observation model, we provide the sample complexity $\tilde{\mathcal{O}}\left(\min\left\{\frac{t_{\text{mix}}}{\epsilon^{2}}\frac{1}{(1-\gamma)^{6}d_{\min}^{3}},\frac{1}{\epsilon}\frac{\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\sigma_{2}(\bm{W}))(1-\gamma)^{4}d_{\min}^{3}}\right\}\right)$ in terms of the infinity norm, which appears to be the first result for sample complexity under tabular setting without any strong assumptions. Furthermore, we also provide a sample complexity result for the independent and identically distributed (i.i.d.) observation model.
@@
    Our analysis relies on switched system modeling of Q-learning, providing new insights for interpretation of distributed Q-learning algorithms.
@@
The non-asymptotic behavior of distributed TD-learning was studied in Doan et al. [2019], Sun et al. [2020], Wang et al. [2020], Lim and Lee [2023], which were motivated from the distributed optimization and control literature [Nedic and Ozdaglar, 2009, Wang and Elia, 2010, Pu and Nedić, 2021]. Distributed versions of various TD-learning algorithms were investigated in Macua et al. [2014], Lee et al. [2018], Wai et al. [2018], Stanković et al. [2023]. As for actor-critic algorithm [Konda and Tsitsiklis, 1999], its extension to distributed setting was studied in Zhang et al. [2018a, b], Zhang and Zavlanos [2019], Suttle et al. [2020], Zeng et al. [2022a]. Moreover, Zhang et al. [2021] investigated distributed algorithm for fitted Q-iteration, which is similar to solving a least squares problem. Meanwhile, a line of research has focused on dealing with exponential scaling in the action space Lin et al. [2021], Qu et al. [2022], Zhang et al. [2023], Gu et al. [2024].
The distributed Q-learning algorithm under the setting when only the local reward is observable, was first studied by Kar et al. [2013]. They proposed the so-called QD-learning proving asymptotic convergence using two-time scale stochastic approximation approaches. Zeng et al. [2022b], Heredia et al. [2020] proved finite-time bounds of distributed Q-learning with linear function approximation. However, the works require additional strong assumptions, which may not hold even in the tabular setup. In particular, Zeng et al. [2022b] considered a strongly monotone condition to hold, and Heredia et al. [2020] posed a particular assumption on the state-action distribution.
@@ 2.1 Multi Agent MDP
A multi-agent Markov decision process (MAMDP) consists of the tuple $({\mathcal{S}},\{{\mathcal{A}}_{i}\}_{i=1}^{N},{\mathcal{P}},\{r^{i}\}^{N}_{i=1})$, where ${\mathcal{S}}:=\{1,2,\dots,|{\mathcal{S}}|\}$ is the finite set of states, ${\mathcal{A}}_{i}:=\{1,2,\dots,|{\mathcal{A}}_{i}|\}$ is the finite set of actions for each agent $i\in{\mathcal{V}}$, ${\mathcal{P}}:{\mathcal{S}}\times\prod^{N}_{i=1}{\mathcal{A}}_{i}\times{\mathcal{S}}\to[0,1]$ is the transition probability, and $r^{i}:{\mathcal{S}}\times\prod^{N}_{i=1}{\mathcal{A}}_{i}\times{\mathcal{S}}\to\mathbb{R}$ is the reward function of agent $i\in{\mathcal{V}}$. We will use the notation ${\mathcal{A}}:=\prod^{N}_{i=1}{\mathcal{A}}_{i}=\{1,2,\dots,|{\mathcal{A}}|\}$ where tuple of actions are mapped to unique integer.
@@
$Q^{*}(s,{\bm{a}})=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}r^{i}(s,{\bm{a}},s^{\prime})+\gamma\max_{{\bm{u}}\in{\mathcal{A}}}Q^{*}(s,{\bm{u}})\right],\quad\forall s\in{\mathcal{S}},{\bm{a}}\in{\mathcal{A}}.$ (1)
Since each agent only has an access to its local reward $r^{i}$, it is impossible to learn the central optimal Q-function without sharing additional information among the agents. However, we assume that there is no central coordinator that can communicate with all the agents. Instead, we will consider a more restricted communication scenario where each agent can share its learning parameter only with a subset of the agents. This communication constraint can be caused by several reasons such as infrastructures, privacy, and spacial topology. The communication structure among the agents can be described by an undirected simple connected graph ${\mathcal{G}}:=({\mathcal{V}},{\mathcal{E}})$, where ${\mathcal{V}}$ denotes the set of vertices and ${\mathcal{E}}\subset{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges. Each agent will be described by a vertex $v\in{\mathcal{V}}:=\{1,2,\dots,N\}$, where $N$ is the number of agents. Moreover, each agent $i\in{\mathcal{V}}$ only communicates with its neighbours, denoted as ${\mathcal{N}}_{i}:=\{j\in{\mathcal{V}}\mid(i,j)\in{\mathcal{E}}\}$.
To further proceed, we will use the following matrix and vector notations:
${\bm{P}}$ $:=\begin{bmatrix}{\bm{P}}_{1,1}&{\bm{P}}_{1,2}&\cdots&{\bm{P}}_{|{\mathcal{S}}|,|{\mathcal{A}}|}\end{bmatrix}^{\top}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|\times|{\mathcal{S}}|},\quad{\bm{R}}^{i}:=\begin{bmatrix}{\bm{R}}^{i\top}_{1}&\cdots&{\bm{R}}^{i\top}_{|{\mathcal{S}}|}\end{bmatrix}^{\top}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|},$
where for $s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}$, ${\bm{P}}_{s,{\bm{a}}}\in\mathbb{R}^{|{\mathcal{S}}|}$ and ${\bm{R}}^{i}_{s}\in\mathbb{R}^{|{\mathcal{A}}|}$ are column vectors such that $[{\bm{P}}_{s,{\bm{a}}}]_{s^{\prime}}={\mathcal{P}}(s,{\bm{a}},s^{\prime})$ for $s^{\prime}\in{\mathcal{S}}$, and $[{\bm{R}}^{i}_{s}]_{{\bm{a}}}=\mathbb{E}\left[r^{i}(s,{\bm{a}},s^{\prime})\mid s,{\bm{a}}\right]$, respectively. Throughout the paper, we will represent a policy in a matrix form. A greedy policy over ${\bm{Q}}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$, which is denoted as $\pi_{{\bm{Q}}}:{\mathcal{S}}\to{\mathcal{A}}$, i.e., $\pi_{{\bm{Q}}}(s)=arg\max_{{\bm{a}}\in{\mathcal{A}}}({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{Q}}$, can be represented as a matrix as follows:
@@
where ${\bm{e}}_{s}$ and ${\bm{e}}_{{\bm{a}}}$ represent the canonical basis vector whose $s$-th and ${\bm{a}}$-th element is only one and others are all zero in $\mathbb{R}^{|{\mathcal{S}}|}$ and $\mathbb{R}^{|{\mathcal{A}}|}$, respectively, and $\otimes$ denotes the Kronecker product. We can prove that ${\bm{P}}{\bm{\Pi}}^{{\bm{Q}}}$ for ${\bm{Q}}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$ represents a transition probability of state-action pairs under policy $\pi$, i.e., $({\bm{e}}_{s^{\prime}}\otimes{\bm{e}}_{{\bm{a}}^{\prime}})^{\top}({\bm{P}}{\bm{\Pi}}^{{\bm{Q}}})({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})={\mathbb{P}}\left[(s_{k+1},{\bm{a}}_{k+1})=(s^{\prime},{\bm{a}}^{\prime})\mid(s_{k},{\bm{a}}_{k})=(s,{\bm{a}}),\pi_{{\bm{Q}}}\right]$ for $s,s^{\prime}\in{\mathcal{S}}$ and ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{A}}$. Now, we can rewrite the Bellman equation in (1) using the matrix notations as follows:
${\bm{R}}^{\mathrm{avg}}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}^{*}}{\bm{Q}}^{*}={\bm{Q}}^{*},$
where ${\bm{R}}^{\mathrm{avg}}=\frac{1}{N}\sum^{N}_{i=1}{\bm{R}}^{i}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$ and ${\bm{Q}}^{*}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$ represents optimal Q-function, $Q^{*}$, i.e., $({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{Q}}^{*}=Q^{*}(s,{\bm{a}})$ for $s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}$.
@@ 2.2 Distributed Q-learning
${\bm{Q}}^{i}_{k+1}(s_{k},{\bm{a}}_{k})=$ $\sum_{j\in{\mathcal{N}}_{i}}[{\bm{W}}]_{ij}{\bm{Q}}^{j}_{k}(s_{k},{\bm{a}}_{k})+\alpha\left(r^{i}_{k+1}+\gamma\max_{{\bm{a}}\in{\mathcal{A}}}{\bm{Q}}^{i}_{k}(s^{\prime}_{k},{\bm{a}})-{\bm{Q}}^{i}_{k}(s_{k},{\bm{a}}_{k})\right)$ (2)
@@
For all $i\in[N]$, $[{\bm{W}}]_{ii}>0$ and $[{\bm{W}}]_{ij}>0$ if $(i,j)\in{\mathcal{E}}$, otherwise $[{\bm{W}}]_{ij}=0$. Furthermore, $\sum_{j=1}^{N}[{\bm{W}}]_{ij}=\sum_{i=1}^{N}[{\bm{W}}]_{ji}=1$, and ${\bm{W}}$ is as symmetric matrix, i.e., ${\bm{W}}^{\top}={\bm{W}}$.
The assumption is widely adopted in the literature of distributed learning scheme [Nedic and Ozdaglar, 2009, Kar et al., 2013, Heredia et al., 2020, Zeng et al., 2022b]. In Appendix B, we provided a simple strategy to construct the doubly stochastic matrix by communicating only with its neighbour.

### v2

@@ Abstract
Multi-agent reinforcement learning (MARL) has witnessed a remarkable surge in interest, fueled by the empirical success achieved in applications of single-agent reinforcement learning (RL). In this study, we consider a distributed Q-learning scenario, wherein a number of agents cooperatively solve a sequential decision making problem without access to the central reward function which is an average of the local rewards. In particular, we study finite-time analysis of a distributed Q-learning algorithm, and provide a new sample complexity result of $\tilde{\mathcal{O}}\left(\max\left\{\frac{1}{\epsilon^{2}}\frac{t_{\text{mix}}}{(1-\gamma)^{6}d_{\min}^{4}},\frac{1}{\epsilon}\frac{\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\sigma_{2}(\bm{W}))(1-\gamma)^{4}d_{\min}^{3}}\right\}\right)$ under tabular lookup setting for Markovian observation model.
@@ 1 Introduction
MARL has been studied under various scenarios including an access to central reward function [Tan, 1993, Claus and Boutilier, 1998, Littman, 2001]. In particular, our interest lies in the the distributed learning paradigm where agents collaborate to solve a shared problem, constrained to communicate solely with their neighboring agents and does not have access to central reward function. Such setting has came of interest due to its wide applications [Blumenkamp et al., 2022, Prabuchandran et al., 2014, Zhao et al., 2021]. Compared to scenarios where a centralized coordinate exists, the distributed paradigm has advantage in terms of privacy-preservation and scalability. One notable example is the distributed adaptation of temporal-difference (TD) learning, as demonstrated in studies by Doan et al. [2019], Wang et al. [2020], Lim and Lee [2023], to name a few.
Meanwhile, in the literature of single-agent RL, Q-learning [Watkins and Dayan, 1992] is one of the most important algorithms in RL. The non-linear max-operator in Q-learning algorithm imposes difficulty in the analysis, and its non-asymptotic analysis has been an active research area recently [Even-Dar et al., 2003, Chen et al., 2021, Lee et al., 2023, Li et al., 2024]. However, distributed learning framework for Q-learning has not been studied in detail. In particular, distributed Q-learning has been studied in an asymptotic sense [Kar et al., 2013], i.e., the algorithm converges over time as it approaches infinity, or in a non-asymptotic sense under additional assumptions on the problem [Heredia et al., 2020, Zeng et al., 2022b]Wang et al. [2022] studied a version of distributed Q-learning in tabular setting but differs from the one in Kar et al. [2013]. This motivates our study to understand its non-asymptotic behavior under tabular setup, i.e., all the state-action values are stored in a table. Our contribution can be summarized as follows:
@@
    For Markovian observation model, we provide the sample complexity $\tilde{\mathcal{O}}\left(\max\left\{\frac{t_{\text{mix}}}{\epsilon^{2}}\frac{1}{(1-\gamma)^{6}d_{\min}^{4}},\frac{1}{\epsilon}\frac{\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\sigma_{2}(\bm{W}))(1-\gamma)^{4}d_{\min}^{3}}\right\}\right)$ in terms of the infinity norm under tabular setting. We derive, for the first time, the finite-time analysis of QD-learning [Kar et al., 2013] in its original form, which is one of the most fundamental and widely used distributed Q-learning methods. While several works have addressed other types of distributed Q-learning, the analysis of QD-learning has remained unexplored until now. Furthermore, we also provide a sample complexity result for the independent and identically distributed (i.i.d.) observation model.
@@
    Our analysis relies on switched system modeling of Q-learning, providing new insights for interpretation of distributed Q-learning algorithms. We show that the distributed Q-learning also allows switched system interpretation as in the single-agent case.
@@
The non-asymptotic behavior of distributed TD-learning was studied in Doan et al. [2019], Sun et al. [2020], Wang et al. [2020], Lim and Lee [2023], which were motivated from the distributed optimization and control literature [Nedic and Ozdaglar, 2009, Wang and Elia, 2010, Pu and Nedić, 2021]. Distributed versions of various TD-learning algorithms were investigated in Macua et al. [2014], Lee et al. [2018]. As for actor-critic algorithm [Konda and Tsitsiklis, 1999], its extension to distributed setting was studied in Zhang et al. [2018a, b], Zhang and Zavlanos [2019], Zeng et al. [2022a]. Meanwhile, Yang et al. [2023] considered a distributed policy gradient approach. Moreover, Zhang et al. [2021] investigated distributed algorithm for fitted Q-iteration, which is similar to solving a least squares problem. Furthermore, a line of research has focused on dealing with exponential scaling in the action space Lin et al. [2021], Qu et al. [2022], Zhang et al. [2023], Gu et al. [2024].
The distributed Q-learning algorithm under the setting when only the local reward is observable, was first studied by Kar et al. [2013]. They proposed the so-called QD-learning proving asymptotic convergence using two-time scale stochastic approximation approaches. Zeng et al. [2022b], Heredia et al. [2020] proved finite-time bounds of distributed Q-learning with linear function approximation. However, the works require additional strong assumptions, which may not hold even in the tabular setup. In particular, Zeng et al. [2022b] considered a strongly monotone condition to hold, and Heredia et al. [2020] posed a particular assumption on the state-action distribution. Wang et al. [2022] studied a distributed Q-learning model motivated from the adapt-then-combine scheme [Chen and Sayed, 2012] in the distributed optimization literature and provided a sample complexity bound in terms of high-probability. The communication process must occur after the update scheme, whereas our model allows the TD-error to be computed concurrently with the communication process.
@@
Meanwhile, a separate line of research focusing on multi-agent problems is the federated reinforcement learning literature [Khodadadian et al., 2022, Woo et al., 2023, Zheng et al., 2023]. This approach differs from the distributed learning scenario in two key aspects: it employs a centralized controller, and all agents share a common reward function.
@@ 2.1 Multi Agent MDP
A multi-agent Markov decision process (MAMDP) consists of the tuple $({\mathcal{S}},\{{\mathcal{A}}_{i}\}_{i=1}^{N},{\mathcal{P}},\{r^{i}\}^{N}_{i=1},\gamma)$, where ${\mathcal{S}}:=\{1,2,\dots,|{\mathcal{S}}|\}$ is the finite set of states, ${\mathcal{A}}_{i}:=\{1,2,\dots,|{\mathcal{A}}_{i}|\}$ is the finite set of actions for each agent $i\in{\mathcal{V}}$, ${\mathcal{P}}:{\mathcal{S}}\times\prod^{N}_{i=1}{\mathcal{A}}_{i}\times{\mathcal{S}}\to[0,1]$ is the transition probability, and $r^{i}:{\mathcal{S}}\times\prod^{N}_{i=1}{\mathcal{A}}_{i}\times{\mathcal{S}}\to\mathbb{R}$ is the reward function of agent $i\in{\mathcal{V}}$. We will use the notation ${\mathcal{A}}:=\prod^{N}_{i=1}{\mathcal{A}}_{i}=\{1,2,\dots,|{\mathcal{A}}|\}$ where tuple of actions are mapped to unique integer. $\gamma\in(0,1)$ is the discount factor.
@@
$Q^{*}(s,{\bm{a}})=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}r^{i}(s,{\bm{a}},s^{\prime})+\gamma\max_{{\bm{u}}\in{\mathcal{A}}}Q^{*}(s^{\prime},{\bm{u}})\right],\quad\forall s\in{\mathcal{S}},{\bm{a}}\in{\mathcal{A}}.$ (1)
Since each agent only has an access to its local reward $r^{i}$, it is impossible to learn the central optimal Q-function without sharing additional information among the agents. However, we assume that there is no central coordinator that can communicate with all the agents. Instead, we will consider a more restricted communication scenario where each agent can share its learning parameter only with a subset of the agents. This communication constraint can be caused by several reasons such as infrastructure, privacy, and spatial topology. The communication structure among the agents can be described by an undirected simple connected graph ${\mathcal{G}}:=({\mathcal{V}},{\mathcal{E}})$, where ${\mathcal{V}}$ denotes the set of vertices and ${\mathcal{E}}\subset{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges. Each agent will be described by a vertex $v\in{\mathcal{V}}:=\{1,2,\dots,N\}$, where $N$ is the number of agents. Moreover, each agent $i\in{\mathcal{V}}$ only communicates with its neighbours, denoted as ${\mathcal{N}}_{i}:=\{j\in{\mathcal{V}}\mid(i,j)\in{\mathcal{E}}\}$.
To further proceed, we will use the following matrix and vector notations: ${\bm{P}}:=\begin{bmatrix}{\bm{P}}_{1,1}&{\bm{P}}_{1,2}&\cdots&{\bm{P}}_{|{\mathcal{S}}|,|{\mathcal{A}}|}\end{bmatrix}^{\top},\quad{\bm{R}}^{i}:=\begin{bmatrix}{\bm{R}}^{i\top}_{1}&\cdots&{\bm{R}}^{i\top}_{|{\mathcal{S}}|}\end{bmatrix}^{\top}$ where ${\bm{P}}_{s,{\bm{a}}}\in\mathbb{R}^{|{\mathcal{S}}|}$ and ${\bm{R}}^{i}_{s}\in\mathbb{R}^{|{\mathcal{A}}|}$ are column vectors such that $[{\bm{P}}_{s,{\bm{a}}}]_{s^{\prime}}={\mathcal{P}}(s,{\bm{a}},s^{\prime})$ for $s^{\prime}\in{\mathcal{S}}$, and $[{\bm{R}}^{i}_{s}]_{{\bm{a}}}=\mathbb{E}\left[r^{i}(s,{\bm{a}},s^{\prime})\mid s,{\bm{a}}\right]$, respectively. We assume that $||r^{i}(s,{\bm{a}},s^{\prime})||_{\infty}\leq R_{\max}$ for some positive real number $R_{\max}.$ Throughout the paper, we will represent a policy in a matrix form. A greedy policy over ${\bm{Q}}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$, which is denoted as $\pi_{{\bm{Q}}}:{\mathcal{S}}\to{\mathcal{A}}$, i.e., $\pi_{{\bm{Q}}}(s)=arg\max_{{\bm{a}}\in{\mathcal{A}}}({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{Q}}$, can be represented as a matrix as follows:
@@
where ${\bm{e}}_{s}$ and ${\bm{e}}_{{\bm{a}}}$ represent the canonical basis vector whose $s$-th and ${\bm{a}}$-th element is only one and others are all zero in $\mathbb{R}^{|{\mathcal{S}}|}$ and $\mathbb{R}^{|{\mathcal{A}}|}$, respectively, and $\otimes$ denotes the Kronecker product. We can prove that ${\bm{P}}{\bm{\Pi}}^{{\bm{Q}}}$ for ${\bm{Q}}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$ represents a transition probability of state-action pairs under policy $\pi$, i.e., $({\bm{e}}_{s^{\prime}}\otimes{\bm{e}}_{{\bm{a}}^{\prime}})^{\top}({\bm{P}}{\bm{\Pi}}^{{\bm{Q}}})({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})={\mathbb{P}}\left[(s_{k+1},{\bm{a}}_{k+1})=(s^{\prime},{\bm{a}}^{\prime})\mid(s_{k},{\bm{a}}_{k})=(s,{\bm{a}}),\pi_{{\bm{Q}}}\right]$ for $s,s^{\prime}\in{\mathcal{S}}$ and ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{A}}$. Now, we can rewrite the Bellman equation in (1) using the matrix notations as follows: ${\bm{R}}^{\mathrm{avg}}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}^{*}}{\bm{Q}}^{*}={\bm{Q}}^{*}$, where ${\bm{R}}^{\mathrm{avg}}=\frac{1}{N}\sum^{N}_{i=1}{\bm{R}}^{i}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$ and ${\bm{Q}}^{*}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$ represents optimal Q-function, $Q^{*}$, i.e., $({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{Q}}^{*}=Q^{*}(s,{\bm{a}})$ for $s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}$.
@@ 2.2 Distributed Q-learning
${\bm{Q}}^{i}_{k+1}(s_{k},{\bm{a}}_{k})=$ $\sum_{j\in{\mathcal{N}}_{i}\cup\{i\}}[{\bm{W}}]_{ij}{\bm{Q}}^{j}_{k}(s_{k},{\bm{a}}_{k})+\alpha\left(r^{i}_{k+1}+\gamma\max_{{\bm{a}}\in{\mathcal{A}}}{\bm{Q}}^{i}_{k}(s^{\prime}_{k},{\bm{a}})-{\bm{Q}}^{i}_{k}(s_{k},{\bm{a}}_{k})\right)$ (2)
${\bm{Q}}^{i}_{k+1}(s,{\bm{a}})=$ $\sum_{j\in{\mathcal{N}}_{i}\cup\{i\}}[{\bm{W}}]_{ij}{\bm{Q}}^{j}_{k}(s,{\bm{a}}),\quad s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}\setminus\{(s_{k},{\bm{a}}_{k})\},$
@@
For all $i\in[N]$, $[{\bm{W}}]_{ii}>0$ and $[{\bm{W}}]_{ij}>0$ if $(i,j)\in{\mathcal{E}}$, otherwise $[{\bm{W}}]_{ij}=0$. Furthermore, $\sum_{j=1}^{N}[{\bm{W}}]_{ij}=\sum_{i=1}^{N}[{\bm{W}}]_{ji}=1$, and ${\bm{W}}^{\top}={\bm{W}}$.
The assumption is widely adopted in the literature of distributed learning scheme [Heredia et al., 2020, Zeng et al., 2022b]. In Appendix B, we provided a simple strategy to construct the doubly stochastic matrix by communicating only with its neighbour.